Foreverangel:

When I put it into a calculator it just tells me if it is true or false 59° 62.8° 64.2° 68°

Foreverangel:

1 attachment
Vocaloid:

another law of sines problem - can you show me how you set it up and what program you're using?

surjithayer:

 sine formula \[\frac{ \sin A }{ a }=\frac{ \sin B }{ b }=\frac{ \sin C }{ c }\]

surjithayer:

put A=Z a=29 B=101 b=32 Taking first two solve for Z